A random variable X follows the uniform distribution with a lower limit of 520 and an upper limit of 620.
a. Calculate the mean and the standard deviation for the distribution. (Round intermediate calculation for standard deviation to 4 decimal places and final answer to 2 decimal places.)


b. What is the probability that X is less than 570? (Round intermediate calculations to at least 4 decimal places and final answer to 2 decimal places.)

Answer by Boreal(15235)

mean is (a+b)/2=570, the midpoint.
variance is (1/12)(b-a)^2=1/12(10000)=833.33
sd is sqrt (V)=28.87